[figure: 9.7 cm]

BAND GAP STABILITY IN KONDO INSULATORS

aA. Ślebarski, ^{b}A. Jezierski, ^{c}A. Zygmunt aInstitute of Physics, University of Silesia, 40–007 Katowice, Poland

bInstitute of Molecular Physics, Polish Academy of Sciences, 60-179 Poznań, Poland cInstitute for Low Temperature and Structure Research, Polish Academy of Sciences,

50-950 Wrocław, Poland

We report on magnetic measurements and electronic structure investigations of CeNiSn and CeRhSb. Both belong to the group of Kondo insulators. The magnetic susceptibility shows the nonmagnetic ground state for these compounds and their alloys. The 3d x-ray photoemission spectroscopy spectra show evidence for the mixed valence state of Ce in CeRhSb alloys, as also seen for CeNiSn, whereas the spectra for the La substituted (Ce,La)NiSn compounds show only evidence for a pure Ce^{3+} ground state. We suggest the presence of Kondo-hole states in (Ce,La)RhSb. The location of the pseudogap in CeRhSb varies with the number of free electron, the valence of Ce, and the f − d hybridization. We discuss the similar crystallographic properties and the closed electronic structures of ZrNiSn-type semi-Heusler alloys and CeNiSn-type Kondo insulators.

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : PLENARY LECTURE

Corresponding author : A. Ślebarski

Address for correspondence : Institute of Physics, University of Silesia, ul. Uniwersytecka 4, 40–007 Katowice, Poland

Email address : slebar@usctoux1.cto.us.edu.pl